OpenStudy (anonymous):

Polynomials: How to do? 9-x^2

jhonyy9 (jhonyy9):

(3-x)(3+x)

OpenStudy (anonymous):

What exactly are you trying to do?

OpenStudy (anonymous):

here u first convert it to a^2 - b^2 so that u can then use the formula a^2 - b^2 = (a+b)(a-b) 9 = 3^2 so 9 -x^2 = 3^2 - x^2 = (3+x)(3-x)

OpenStudy (anonymous):

Thanks, How did u get the 3^2?

jhonyy9 (jhonyy9):

x_1,_2 = +/- +

jhonyy9 (jhonyy9):

pardon will b3 +/- 3

jhonyy9 (jhonyy9):

x_1,_2 = +/- 3

OpenStudy (anonymous):

teejackson 9 = 3x3 = 3^2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

now is it clear??

OpenStudy (anonymous):

numbers like 4, 9,16,25 are square numbers 4 = 2x2 = 2^2 9 = 3x3 = 3^2 16 = 4x4 = 4^2 etc.

OpenStudy (anonymous):

Harkirat, no not really this is very hard for me. So (3+x)(3-x) is the formula, but why is the 3 used for both side and not the x^2

OpenStudy (anonymous):

see 9 - x^2 = 3^2 - x^2 = (3+x)(3-x) we are using 3 as well as x in both brackets

OpenStudy (anonymous):

so we take out the ^2 and only worry about the 3"s, because there 2 of them?

OpenStudy (anonymous):

The generic rule is:\[a^2-b^2=(a+b)(a-b)\]So we're trying to turn your equation into that form:\[9-x^2 = 3^2 - x^2\]Now we just substitute "3" for "a" and "x" for "b" in the original equation:\[3^2-x^2=(3+x)(3-x)\]The little 2's mean "squared," which just means the number times itself. The trick here was in recognizing that 9 is 3 squared. 3 x 3 = 9. The term is "square" because if you had a square with sides of length 3, the area would be 9.